Early Sanskritization. Origins and Development of the Kuru State.

Michael Witzel

Summary

The Mahābhārata is, by and large, the tale of a great battle between two sections of the ancient Kuru people of the Haryana/W. Uttar Pradesh region, as told at the snake sacrifice of the post-battle Kuru king Parīks.it. Since the Vedic texts are “tape recordings” of the Vedic period we can use them as basis for judging the less strict tradition of the Epics. To provide such a basis, the present paper is an investigation of the forces behind the formation of the Kuru realm at the end of the R̊gvedic period.

The emergence of the Kuru realm is of extraordinary importance as its civilization has influenced later Indian ritual, society and political formations, frequently even until today. Comments are welcome and will be published in a separate column of letters/discussion.

Kuruks.etra,^{1} the sacred land of Manu where even the gods perform their sacrifices, is the area between the two small rivers Sarsuti and Chautang,^{2} situated about a hundred miles north-west of Delhi. It is here that the Mahābhārata battle took place.^{3} Why has Kuruks.etra been regarded so highly ever since the early Vedic period? Actually, the R̊gvedic archetype of the Mahābhārata, the “Ten Kings’ Battle” (dāśarājña), took place further west on the Parus.n.ī (Ravī). Due to the victory of the Bharata chieftain Sudās in this battle, the Bharata tribe was able to settle in the Kuruks.etra area.^{4} The evolvement of the small tribal Bharata domination into that of a much larger Kuru realm is not recorded by our texts. The Kurus suddenly appear on the scene in the post-R̊gvedic texts. As so often, the Sanskrit texts record only the results of certain developments.

1This article is a summary of my forthcoming monograph: The Realm of the Kuru. Here, I make liberal use especially of the introduction and the final chapters of the monograph and present the results rather than the line of the argument of the underlying investigations. — This paper was first presented at the yearly Japanese Conference on South Asian Studies at Kyoto in October 1989 and a summary was published: The Realm of the Kurus: Origins and Development of the First State in India. Nihon Minami Ajia Gakkai Zenkoku Taikai, Hookoku Yooshi, [Summaries of the Congress of the Japanese Association for South Asian Studies], Kyoto 1989. Subsequently, it was read at the Conference on Indian History, organized by B. Kölver at Munich in June 1992, and elsewhere as well.

2Sarsuti (Ved. Sarasvatī) = Ghagghar, Chautang = Ved. Dr_{}˚s.advatī, in the Thanesar/Sirsa/Hanumangarh area; see H. Wilhelmy, Das Urstromtal am Ostrand der Indusebene und das Sarasvati-Problem. Zeitschrift f. Geomorphologie, N.F. Suppl. Bd. 8, 1969, Yash Pal et al., Remote sensing of the ‘lost’ Saraswati River, in: B.B. Lal, and S.P. Gupta, Frontiers of the Indus Civilisation, Delhi 1984, 499–504.

3Actually, another big battle is attested in the R_{}˚ V, at 1.53.9, one of twenty kings. — Kuruks.etra is well known from various Vedic texts as the offering ground of the gods (devayajana) and from later sources such as the Manu Smr_{}˚ ti, Mahābhārata, Vāmana Purān.a 23.13-40; even today it is visited by many pilgrims.

4See R_{}˚ V 3.53, with Sudās settling in the center, on the Sarasvatī, and the areas conquered east, west and north of it, while the south is expressis verbis excluded as the non-Indo-Aryan land of the Kīkat.a and of Pramaganda.

Electronic Journal of Vedic Studies (EJVS) 1-4 (1995) pp.(1–26)

Michael Witzel

I. THE MIDDLE VEDIC PERIOD AND THE MANTRA EPOCH

This “gap” between the R̊gveda and the other Vedic texts is one of the major dark periods of Indian history; in fact, it often is not even recognized as a separate period by the very scholars who deal with the Vedic texts. However, in my opinion, it is this period (together with the slightly earlier formation of the Bharata realm), which is of crucial importance for the development of all later Indian culture and civilization. It is at this moment that the social “raw material” present in R̊gvedic time was intentionally transformed into what became the core and the pattern first of Vedic and, later on, of Hindu culture.

What we have of this time are only the Vedic texts.^{5} Archaeology has recently, and increasingly so, become another factor in describing this period, although the correlation between the texts and the archaeological facts is still a matter of much discussion. We have to rely on the texts, their form, their organisation, and their language in its historical development as well as in its synchronic dialectic spread, their internal chronology; furthermore we have to take into account the criteria such as that of text formation, that of the development and spread of the various Vedic schools of ritual interpretation, of the development of ritual and religious thought in general, and, of course, the occasional remarks in the texts about the tribes and peoples of the area, their history, their material culture and its gradual development in time, etc.

In doing so, it is notable that apparently small observations, such as those on phonetic peculiarities, on intentional use of “high” R̊gvedic forms in one of the two AV texts (Paippalāda Sam .hitā), the intentional use of outdated, archaic materials in ritual, the taking over of ancient materials into the final R̊V collection and into the YV texts, or the archaization of Atharvaveda Mantras by starting hymns with hieratic meters, help to create a framework for judging the historical developments and the trends in this early culture.

Early Sanskritization

In this procedure, special attention must be paid to the historical levels in the development of the texts, — and not just to their order in Indian tradition: one usually distinguishes Sam .hitās, Brāhman.as, Āran.yakas, Upanis.ads, and Sūtras, in roughly that chronological order. The internal chronology of the texts helps to establish historical levels. Even more so, the development of the Vedic language is a secure guide in doing so: we have to distinguish five text layers^{6} which do not always coincide with the traditional division given just now. These five linguistic and textual levels can conveniently be divided into three major periods which are distinct in language, habitat, and in their social, religious, and political features: the Old Vedic period (level 1: R̊gveda), the Middle Vedic period (levels 2–4a) and the Late Vedic period (levels 4b, 5: the later Brāhman.as, Āran.yakas, Upanis.ads, and most Sūtras).

5For a summary of the texts and the availabe tools for their study as well as the major work done on them, see S. Jamison and M. Witzel, Vedic Hinduism, in A. Sharma, The study of Hinduism (forthc.).

6They are:

1. R_{}˚ gveda (with a late addition, book 10, and also including parts of book 1);

2. Mantra language (Atharvaveda, Sāmaveda, R_{}˚ gvedakhila, the Mantras of the Yajurveda, i.e. MS, KS/KpS,

TS, VS, etc.); 3. Expository prose of the Yajurveda Sam .hitā texts (MS, KS/KpS, TS), 4. the Brāhman.a prose (including the older portions of the Āran.yaka and Upanis.ads, as well as the earliest

Sūtras, such as BŚS, VādhS); 5. the late Vedic Sūtras (and the post-Vedic Upanis.ads).

However, when the Vedic texts are discussed by Vedic and other scholars, they usually are treated as poetry, as ritual handbooks or as early philosophy, that is — only as texts.^{7} Even after some 150 years of study, the Vedic period as a whole does not seem to have a history, and its texts are generally thought to have been composed in a geographical vacuum “somewhere in Northern India”.^{8}

Against this vague background it is perhaps not surprising that the professional writers on older Indian History did not shed much light on the early and middle Vedic period until a few years ago. The communis opinio still is that the R̊V represents a fight of “everybody against everybody else”.^{9} It is only in the recent book on Indian history by H. Kulke and D. Rothermund^{10} that the Vedic period is treated more adequately. In this work, recent progress both in archaeology and in Vedic studies has been made use of and an up-to-date, fairly detailed and quite reliable picture of the period emerges. However, in this paper, I propose to add some sigfnificant features to the evolving picture.

***

The history of the earlier Vedic period can be summarized as follows. The first fixed dates in Indian history that are usually mentioned are that of the Buddha around 500 B.C.^{11} or rather 400 B.C. and that of Pān.ini. Both dates, in fact, presuppose the evolvement of the bulk of Vedic literature. The beginning of the Vedic period, however, is equally vague and uncertain. Recent findings in archaeology, however, put the disintegration of the Indus civilization at c. 1900 B.C. As the R̊V does not speak of cities but only of ruins (armaka),^{12} even larger ones ([mahā-]vailasthāna), we may suppose that the Indo-Aryans immigrated,^{13} or rather, gradually trickled in,^{14} tribe by

7With a few notable exceptions such as those of W. Rau, K. Mylius, M. Sparreboom.

8Cf. Author, Tracing the Vedic Dialects, in Colette Caillat, Dialectes dans les littératures indo-aryennes. Actes du Colloque International..., 16–18 Septembre 1986. Paris (Collège de France, Institut de Civilisation Indienne) 1989, pp. 97–264.

9R. Thapar, A History of India, vol. 1, Harmondsworth (Penguin), 1966 repr. 1979.

10H. Kulke and D. Rothermund, A History of India, New York 1986.

11Or, as H. Bechert now tells us, maybe 100 years later: H.Bechert, “The date of the Buddha reconsidered,” Indologica Taurinensia 10: 29–36; now also: The dating of the historical Buddha / Die Datierung des historischen Buddha, part 1, ed. H. Bechert, Göttingen 1991–2. Cf. now G. Erdosy, The archaeology of early Buddhism, in: N.K. Wagle, Studies on Buddhism in honour of A.K. Warder, Toronto 1993.

12See the discussion by W. Rau, Zur vedischen Altertumskunde, Akad. Mainz, Wiesbaden 1983, p. 36–38.

13Cf. G. Erdosy, Ethnicity in the Rigveda and its Bearing on the Question of Indo-European Origins, South Asian Studies 5, Cambridge 1977: 40; cf. also G.Erdosy, The Prelude to Urbanisation: Ethnicity and the Rise of Late Vedic Chiefdoms, in Early Historic India, ed. R. Allchin, Cambridge, forthc.; A. Parpola, The coming of the Aryans to Iran and India and the cultural and ethnic identity of the Dāsas, Studia Orientalia 64, Helsinki, 1988: 195–302.

— On the currently fashionable denial of any immigration, see J. C. Shaffer, The Indo-Aryan invasions: Cultural myth and archaeological reality in: The people of South Asia, ed. J. R. Lukacz, New York 1984, pp. 77–90; more zealously: Arun Kumar Biswas, The Aryan Myth, in the recent seminar proceedings (Calcutta 1988): Historical Archaeology of India. A dialogue between archaeologists and historians, ed. A Amita Ray and Samir Mukherjee, New Delhi 1990, p. 29–47.

14Cf. G. Erdosy, op. cit.

tribe^{15} and clan by clan, after 1900 B.C.^{16} As a possible date ad quem for the R̊V one usually adduces the Hittite-Mitanni agreement of the middle of the 14th cent. B.C. which mentions four of the major R̊gvedic gods: Mitra, Varun.a, Indra and the Nāsatya (Aśvin).^{17} The next major archaeological date available is that of the introduction of iron^{18} at c. 1200 B.C. It is first mentioned in the second oldest text, the Atharvaveda, as ‘black metal’ (kr̊s.n.a ayas, śyāma ayas) while the R̊V only knows of ayas itself “copper/bronze”.^{19}

Of the three periods in Vedic history mentioned above, the Old Vedic (R̊gveda) and the Late Vedic periods (Brāhman.as, Upanis.ads, etc.) differ from each other in many respects. It is necessary, first, to characterize the Old and the Late Vedic period briefly.

THE OLD VEDIC PERIOD:  THE LATE VEDIC PERIOD:
R̊gveda, the oldest text  Late Brāhman.as/Early Upanis.ads

geographical area:

Afghanistan, Panjab and surroundings up to the Yamunā (once, the Gaṅgā); all of Northern India, from the Kabul river (Gandhāra) to Aṅga, Pun.d.ra (Bengal), and to Vidarbha (N. E. Mahārās.t.ra), Andhra in the south

political set-up:

some 50 smaller tribes, in constant conflict (gavis.t.i) against each other and against some the aboriginees (dasyu) The Vedic tribes are sometimes arranged into 5 “peoples” (kr̊s.t.i, jana), etc.: 4 in the 4 directions, with major tribe at the “center” two major groups, the Kuru-Pañcāla and Kosala-Videha; at the borders of these units there are some minor tribes: Matsya, Uśīnara, etc.; the area is divided into some 16 “kingdoms”; the Kuru-Pañcāla form the center, the minor tribes and “outsiders” (Bāhīka, Magadha etc.) constitute the outward frames society:

15In my opinion, the earlier ones of the Yadu-Turvaśa, Anu-Druhyu tribes, and later ones such as the combined Pūru-Bharata, who split into two groups upon their arrival in the Afghani borderlands. The Bharatas were the last to move eastwards into the Panjab and into Kuruks.etra and this is represented especially by R_{}˚ V 3 and 7. See Author, “Rigvedic history: poets, chieftains and polities,” in G. Erdosy (ed.), The Indo-Aryans of Ancient South Asia. Language, Material Culture and Ethnicity, (Indian Philology and South Asian Studies, Vol.1, ed. A. Wezler and M. Witzel), Berlin-New York 1995.

16Some overlap of the immigrating Indo-Aryans with the later stages of the Indus civilization is, of course, possible, but should be demonstrated.

17The Mitanni had been exposed to early Indo-Aryan (not: Indo-Iranian) influences a few hundred years earlier, exerted by a branch of those tribes who entered the Bactro-Margiana area around 2100 B.C. and who then proceeded to India. See P. Thieme, Kleine Schriften, Wiesbaden 1971, 396–412; cf. M. Mayrhofer, Die Arier im vorderen Orient — ein Mythos? Wien 1974.

18Apparently from central India, not from Iran!

19See W. Rau, Metalle und Metallgeräte im vedischen Indien, Akad. Mainz (Wiesbaden: 1973); Zur vedischen Altertumskunde, Akad. Mainz, Wiesbaden 1983. AV 11.3.7, 9.5.4, PS 16.53.12

chieftains (rājan) lord over fellow rājanya/ks.atriya (nobility) and the viś “the people”, with the addition of the aboriginees and servants / slaves (dāsa, dasyu, purus.a) front of the ks.atriya and the Brahmins (brahma-ks.atra) against the “people” (viś); successively stricter stratification into the 3 ārya (twice-born) and the additional śūdra (aboriginal) classes (varn.a)

texts and ritual:

gods are invited to often elaborate rituals, such as the soma ritual, they are treated as guests, fed and praised by poets who are inspired and compose hymns in the traditional (IIr./I.E.) the poetical language and traditional meters; the hymns are collected in small sets by the poets’ families and clans the ritual has been transformed into quite an elaborate framework of complicated, framelike structures, set according to two major patterns (soma, is.t.i); poetry of ancient style is no longer produced; most older poetry is collected in some major texts and used in in the ritual in a rather schematic way. New forms of literature dealing with the explanation of the ritual have developed.

The intervening period, i.e. the Middle Vedic epoch, is represented by the Mantras and the expository prose of the YV Sam .hitās (MS, KS/ KpS, TS) and by several older Brāhman.as^{20} — texts composed in the Kuru-Pañcāla area, between Eastern Panjab and Kausambi/Allahabad. The geographical center of Vedic civilization thus has spread from the the Gandhāra/Panjab area to the Eastern border of the Panjab (Kuruks.etra, Haryana) and beyond, well into Uttar Pradesh. Both Sam .hitās of the AV attest the borders of geographical knowledge of this period: they are Balhika (Bactria), and Gandhāri in the north-west while the south-east is marked by the Kāśi (PS) viz. Aṅga (in the somewhat later ŚS).^{21}

II. EMERGENCE

However, the origin of the new large Kuru tribe is still unclear: earlier tribes were remembered as forming parts of the new tribal union, such as the Krivi among the Pañcāla.^{22} In fact the great chieftain of the Kuru still is called chief (rājan) of the Bharata.^{23} In addition, some very neglected passages in Middle Vedic texts suggest that even among the Kuru “dominion is threefold”^{24} and it was six-fold^{25} (originally threefold as well)^{26} among the Pañcāla, which may suggest phyle/tribuslike divisions of these larger unions.^{27} Both tribes, the Kurus and the Pañcālas, form a “people”, of two large “tribes” with separate chieftains whose families, however, intermarry.^{28} In other respects as well, the two tribes form a ritual union within a large chiefdom; it is based on competition between two moieties: for example, they exchange their roving bands of vrātyas (see below). Most interestingly, the pottery of the period seems to echo the tribal differences between the Kurus and Pañcālas and it remains to be seen whether further distinguishing archaeological traits can be identified.^{29}

20The older portions (1–5) of the Aitareya Br.; TB; the lost Śāt.yāyan.a Br. which was elaborated as JB; the older, lost form of PB (pace Bodewitz, JB 1.66–364, introd. p. 2 sq.); some older portions of the largely lost Kat.hB.

21Only occasionally the Vaideha, Saindhava horses and cows (see Localisation p. 181: KS 13.4:183.17, MS 2.5.3:50.10; TS 2.1.4.4, cf. p.183, 195 n. 76;), or the Himalayan mountains are mentioned.

22Vedic Index I,198; Kraivya Pañcāla ŚB 13.5.4.7.

23In the Mantra collection of the royal consecration in Taittirīya Sam .hitā: TS 1.8.10.2, TB 1.7.4.2, 6.7 es.a vo Bharatā rājā; MS 2.6.9:69.7, KS 15.7:214.1 are vague: es.a te janate rājā; VSK 11.3.3, 6.3 es.a vah. kuravo rājā; VSM 9.40 es.a vo ’mī rājā; cf . Keith, TS transl., p. xciii, Author, Localisation, esp. p. 177 sqq. and 182, n. 42.

We now know that the linguistically defined period of the Mantra language^{30} (level 2) intervened between the R̊V (level 1) and the beginning of the Middle Vedic, which is first attested as the expository prose in the “Brāhman.a style” (level 3) of the earliest extant YV Sam .hitās.^{31}

This dark age, the “gap” between the late R̊V and the Mantras of the early YV Sam .hitās, can be approached by asking such questions as: what was the reason for the shift in the geographical location of the tribes from the Panjab to Kuruks.etra and Pañcāla; for the shift of the political center; for the disppearance or unification of the 50-odd major clans and tribes into a few large tribes; for the importance of Kuruks.etra in general; for the development of the the new Vedic (Śrauta) ritual, such as the new order of priests, multiplication of ritual fires, development of new rituals, such as the agnicayana ritual; for certain changes in religion: development of new gods such as Prajāpati, beginning already in R̊V 10; for the collection of the R̊gvedic hymns and other texts; for the differences in language and order of the texts as preserved by different schools of the same Veda: AVŚ : PS, KS : MS, TS; JS : KauthSV?

The Mantra period proper can be characterized as the time of the establishment of the Kuru realm. One or more persons had the ingenious idea to use whatever was present and prominent in the religion and society at the time and to reshape and tailor all these elements in order to establish and maintain Bharata/Kaurava and Kuru dominance.^{32} As will be detailed below, this affected and involved traditional ritual, the institutions of priests, including their number and character, their traditional poetry and ritual texts; furthermore, whatever was amenable to change in the other tribal elites, such as the families of the high aristocracy and the gentry, the poets and bards, and even the leadership of the settled aboriginal population (Nis.āda).

24The rās.t.ra: Vaitahavya, Mitravat, JB 3.196: par. 196; the third group most probably is that of the reigning clan, the Bharata. — Note that this kind of division is still reflected in the Mahābhārata, with two Kuru groups, the Pān.d.ava and Kaurava, and with their two “capitals” at Indraprastha and Hastinapura.

25It is significant that the YV school of the Pañcāla, the Taittirīya, has 6 subschools as well (Baudh., Vādh., Bhār., Āp., Hir., Vaikh.) which echo, also in location (see Author, Localisation, p. 205), the division of this great tribe.

26KS 30.2:183,17, speaks only of a threefold division JB 3.156 tvatvādr_{}˚śās. s.ad. rājānah. Pañcāles.u vedyā iti.

(Rau, Staat und Gesellschaft im alten Indien, nach den Brāhman .as dargestellt, Wiesbaden 1957, p. 47: “Es gibt bekanntlich sechs wie du [Abhipratāran.a] hochadlige Männer”.

27And which provide a parallel to the much later development of Greek city states and the early Roman kingdom and republic, see G. Nagy, Greek mythology and poetics, Ithaca (Cornell Univ. Press), p. 276 sqq.

28See intermarriage at JB 2.278–9.

29See G. Erdosy, Urbanisation in Early Historic India. Oxford 1988; Erdosy, The Prelude to Urbanisation, forthc.

— Note that there is at least one central settlement (“capital”) with a brick building.

30See J. Narten, Die Sprache 14, 113–134, and Author, Tracing, p. 124.

31The texts concerned are the two Atharvaveda texts: Paippalāda and Śaunaka Sam .hitā; the R_{}˚ gveda Khilas (Scheftelowitz’ Apokryphen); the Sāmaveda Sam .hitās: Kauthuma/Rān.āyan.īya Sam .hitā (SV) and Jaiminīya Sam .- hitā (JS), as far as they actually differ from their direct source, the R_{}˚ gveda; and finally the Mantras of the Yajurveda Sam .hitās of the Maitrāyan.ī (MS), Kat.ha (KS), Kapis.t.hala (KpS), Taittirīya (TS), and the Vājasaneyi (VS) schools:

the Kān.va (VSK) and the Mādhyandina Sam .hitās (VSM).

When and where did this take place and who were the main actor(s)? The clue to the enigma is traceable by an investigation into the Kuntāpa ritual (R̊VKh 5, ŚS 20.127). The Kuntāpa section of the R̊gveda Khilas is a very enigmatic but intriguing small collection of hymns and a few prose Mantras (yaju\begin{matrix}s \\ . \end{matrix}). It forms part of the mahāvrata day, i.e. the culmination point of the one-year gavām ayana rite at winter solstice. The main idea seems to be that of helping the sun around its ‘turning point’ at winter solstice. The procedure is assisted by sympathetic magic, such as chariot races imitating the elliptic course of the sun around its turning point. But the Kuntāpa rite also is a fertility rite^{33} and some of the hymns have a curious relation to royal fame and power.^{34}

The name Kuru occurs first as part of the name of a person in the late R̊V,^{35} and then, independently, in the Kuntāpa section as the Kaurava clan/tribe (Kaurama^{36}), where the reign of one of their chiefs is described^{37} as the golden age of the Kaurava / Kuru people under their Great Chief (Kaurava, Ruśama, cf. Kauravya pati). The verses themselves tell us when they were composed: their language is that of the Mantra period. This important yearly ritual transports us into the center of early Kuru power, to Kuruks.etra.^{38} In these stanzas, the ritual is mentioned as taking place with the +Kaurava (Kaurama) among the Ruśama, in Kuru territory. At 5.10.2 a member of this tribe is called a Kauravyah. patih.. His king’s reign apparently constitutes the high point in the history of the tribe. It is clearly described as such:^{39}

“Listen to the good praise of the King belonging to all people, who, (like) a god, is above men, (listen to the the praise) of Pariks.it! - ‘Pariks.it has just now made us peaceful dwelling;^{40} darkness has just now run to its dwelling.’ The Kuru householder, preparing (grains) for milling, speaks (thus) with his wife. — ‘What shall I bring you, sour milk, the mantha [a barley/milk drink], (or) the Parisrut [liquor]?’ the wife keeps asking in the Realm of King Pariks.it. — By itself, the ripe barley bends heavily (iva) over the deep track of the path. The tribe thrives auspiciously in the Realm of King Pariks.it.”^{41}

32Note that the name of the Kurus is a new one (cf., however, Old Persian Kuruš / Cyrus); apparently it was originally a nickname (K. Hoffmann in KEWA III 677); the Kuru kings typically have names that include the denigrating elements duh.-, ku-, such as dur-yodhana, duh.-śāsana; (R_{}˚ V dur-gaha is probably only the horse of a Pūru king, see H.-P. Schmidt, Fs. Heesterman), ŚB 13.5.4 Bharata Dauh.-śānti; Dus.-t.arītu Paum .sāyana, a Kauravya king, ŚB 13.9.3.2 Uccaih.śravas, a Kuru king, the son of Ku-yava; the Pañcāla king Dur-mukha Pañcāla AB 8.14/8.19.

33The mahānagnikā prostitute, and a brahmacārin tease and challenge each other; she and a māgadha man copulate in a hut on the offering ground.

34As does PS 10, which already mentions a Śrauta sava rite, and PS 18.15-26 = ŚS 13).

35Kuruśravan.a Trāsadasyava and the Kān.va poet Kuru-suti (appearing in the R_{}˚ V Anukraman.ī).

36See K. Hoffmann, AzI p. 1 sqq.

37N.B. in the present tense! This indicates contemporaneousness of the author of the hymn and of the king.

38Cf. that the early Krivi, who later make out part of the Pañcālas, defeat, at 8.51.8–9, the Ruśama (several times in R_{}˚ V), and a Ruśama chieftain R_{}˚n.am .caya at 5.30.12–14. This would fit the usual pattern of Pūru, Krivi > Pañcāla, Iks.vāku, and Bharata, Ruśama , etc. > Kuru.

39Bloomfield, Hymns of the Atharva-Veda. (Sacred Books of the East 42.) Oxford 1897, repr. Delhi 1964, already used the same words, but he takes vaiśvanārá- pariks.ít-, following AB 6.32.10, cf. KB 30.5, GB 2.6.12 ŚŚS 12.17, as a name of Agni. Note the more popular Anus.t.ubh meter (for which see H. Oldenberg, Kleine Schriften, 1188 sqq.)

40This is a pun on Pariks.it and Agni. Pariks.it as epithet of Agni means ‘casting light all around’, cf. the use of the word (of Heaven-Earth) in R_{}˚ V 1.123.7, 3.7.1, 10.65.8.

The hymn sums up the good life of this period: peaceful settlement (ks.ema), not strife and war; a variety of food and drink: barley flour, sour milk, the mixture of barley and milk (mantha), a sort of herbal alcohol (parisrut), and a rich harvest of barley.^{42}

Even the exact timeframe is indicated: after Sudās’ Ten Kings’ Battle, which is mentioned at R̊VKh 5.14.1 as dāśarājñe’ mānus.am, the mānus.a (locality)^{43} at the Ten Kings’ Battle. The language of the stanzas affirms this date.^{44} The R̊gvedic social institution of vidatha (5.12.1 vidathyà)^{45} is still known: The Kuru king must have regularly distributed the booty of raids and wars. The most important point, however, is the early post-R̊gvedic praise of the golden age of the Kurus under their King Pariks.it, the ancestor of the well-known Janamejaya Pāriks.ita of Brāhman.a and Mahābhārata fame and of the Pāriks.ita dynasty of the Kurus.

The most important political result of the reform carried out by the dynasty of Pariks.it was the formation of the Kuru tribe and the permanent establishment of the Bhārata-Kuru chiefdom. The formation of the Kuru state and the establishment of its new socio-religious basis is a lasting feature of the Vedic period, and not a transient one like that of the Pūru or Bharata realms in R̊gvedic times. In fact, as we shall see, the “new order” has its distant effects until today.

III. STRATEGIES

The changes were carried out in the center of political power and of contemporary culture, in Kuruks.etra, which now also became the center of the newly emerging Vedic orthopraxy and “orthodoxy”.^{46} At this time, various R̊gvedic tribes in eastern Panjab , Haryana and western Uttar Pradesh first fused into one tribe, the Kuru, which later expanded into two major tribes, the Kuru-Pañcāla. The Kuru union and the realm of their Bhārata/Kaurava kings represents the first larger polity or “early state”^{47} on Indian soil. Its sheer size among the few other surviving R̊gvedic tribes would have insured their dominance.

41r´ājño viśvajan´īyasya yó devó mártā ˙m̆ áti vaiśvanārásya sus.t.utím ´ ā ^{+}śr_{}˚ notā Pariks.ítah.

“Pariks.ín nah. ks.émam akarat táma āsanám ´ ā ^{+}saran”

+marāyyáṅ kúrvan Káuravyah. pátir vadati jāyáya “katarát ta ´ ā harān.i dádhi mánthām3 parisrútam”?

jay´ā pátim ví pr_{}˚ cchati rās.t.ré R´ājñah. Pariks.ítah.

ábhīva svàh. prá jihīte yávah. pakváh. pátho bílam jánas sá bhadrám edhate rās.t.ré R´ājñah. Pariks.ítah.

42Barley fits the Kuruks.etra area; rice is found further eastwards at the time (where wild rice is indigenous), though some rice has been used already in the late Indus civilization. See I. C. Glover, South Asian Archaeology 1977, 7–37.

43H.-P. Schmidt (Indica, Vol. 17, 1980, pp. 41–47) takes mānus.a at R_{}˚ V 7.18 not as a locality (as it clearly is in JB etc.) but as meaning “human world”; cf. also Author, Eastern Iran and the Atharvaveda. Persica 9, 86–128; for mānus.a cf. also EWA II 309.

44The injunctive is still used while it occurs just a few dozen times in AV: likhat; the R_{}˚ gvedic particle gha (5.15.3–4), devatta- (5.15.8a), akr_{}˚n .oh. instead of later (AV) akaro\begin{matrix}h \\ . \end{matrix} (5.21.2).

45See Kuiper, The ancient Indian verbal contest, IIJ 4, 1960, 217–281.

46This is also evident if we trace the movements and differences of the various Vedic schools backwards: The East (Kosala/Videha) has the later schools (Vājasaneyin); the Central area (U.P.: Taittirīya) shows a clear dependence on the western (Kuru) KS/ MS traditions; the South (M.P. north of the Vindhyas: Jaiminīya etc.) is equally dependent on the earlier Central (Pañcāla) schools; there remains, thus, the Kuru territory with schools such as the Maitrāyan.īya and Kat.ha as the nucleus.

The Kuru realm matches many of the characteristics of early states which emerged from traditional tribal chiefdoms or from the larger aggregation of such chiefdoms.^{48} The new Kuru king, in fact, may often still be characterized as a Great Chief.^{49} He is only the primus inter pares (śres.t.hah. svānām) among the high nobility of the Kuru confederation which is characterized as having three subtribes.^{50} However, the new powerful kingship is at least semi-hereditary,^{51} bolstered by a claim of ultimately divine ancestry^{52} re-enacted in ritual. This claim is supported by the royal priest (purohita) and by a retinue of ratnins, royal officials who are bound to the Kuru lord by loyalty and liberal gifts. The new order is further sustained by some major changes in society, such the incipient stratification into four “classes” (varn.a), first met with in the late R̊V (10.90), the establishment of the new priestly corporations representing the Four Vedas, and especially by the ever-increasing dominance of the Brāhma-Ks.atra alliance.^{53} It was created, as some Brāhman.a texts clearly say with Marxist analysis before its day, in order to exploit the rest of the population.^{54}

The establishment of the new Kuru order differed qualitatively from the more gradual R̊gvedic political and social developments. In R̊gvedic times, there clearly were some “non-Āryan”^{55} chiefs such as Varo Suśāman, Balbūtha, Br̊bu^{56} who followed Indo-Aryan religion. They represent examples of an early wave of acculturation.^{57} Even the hieratic poetic language of the R̊V hymns

47Note also that the Indus civilization had about 5 centers i.e. large cities, including one in Gujarat, according to our present state of knowledge: that means no longer “dual capital cities.”

48Cf. H.J.M. Claessen, Current Anthropology Aug.–Oct. 1984, 365–379; H.J.M. Claessen and P. Skalnik, The Early State. The Hague, 1978.

49I will use “king” to designate the ruler of the new Kuru tribe as to underline his preeminence in comparison with the chiefs of the Matsya, Uśīnara, Satvant, etc. tribes who survived from the R_{}˚ gvedic period as smaller independant units.

50Bharata/Tr_{}˚ tsu, Vaitahavya, Mitravat; cf. above n. 23.

51In the R_{}˚ V, and later on, it is clear that kings were be elected, but at the same time, there existed dynasty-like lineages: obviously, the chieftains could be elected from a larger group of noblemen (note that R_{}˚ V 10.90 only speaks of Rājanyas, not of Ks.atriyas!) In one case, that of dus.t.arītu (ŚB 12.9.3.2), we hear of an uninterrupted succession in ten generations before he was ousted. Note that a ritual such as the Rājasūya was necessary to keep the line of the King intact, in case he had no direct heir: by this ritual, he could adopt one (see H. Falk, Die Legende von Śunah.śepa vor ihrem rituellen Hintergrund, ZDMG 134, 1984, 115–135).

52For Rohita, see AV 13 = PS 18.15–26 = AV 13, and his relation to the sun, varcas (x^{v}ar^{e}nah), his identification with Indra, the king of the gods etc. — cf. See Tsuchiyama, Veda.no várcas, Indotetsugaku Bukkyogaku 5, 1990, 67–80.

53Note that the Brahmins pretend to be pre-eminent and semi-independent: “Soma is our king” they say in the royal consecration. The texts, however, also stress that the nobility is the “eater” and the Brahmins are their “food”, see Rau, Staat, p.34 n.6.

54See Rau, Staat, p. 118, A. Weber, Indische Studien 10, 26-35.

55See now also Kuiper, Aryans in the R_{}˚ gveda, Amsterdam-Atlanta 1991.

56Br_{}˚ bu, “lording over the Pan.is on the Gaṅgā”, 6.45.31 (part of a late addition to the hymn). His grandfather had immigrated; Br_{}˚ bu fights with the Bharatas.

57This must have been fairly wide-spread and thorough as not just the names of kings but even the linguistic features of Vedic Sanskrit indicate acculturation.

shows the increasing influence of the local substrate.^{58}

Now, under the Kuru kings, acculturation was followed by well-planned Sanskritization^{59} representing major changes in social format.^{60} It included, in a strategically advantageous way for the Kuru, the older (R̊gvedic) elements of ritual with its priests, texts, and language, while exceedingly stressing its traditional character by being overly archaic^{61} and restrictive. The new class (“caste”) system introduced non-Aryans such as the Śūdras into the Vedic society^{62} but, at the same time, barred them from ritual (and thus, from heaven). Only by way of exception, prominent non-Aryans such as the nis.āda-sthapati and a “border line” artisan, the rathakāra,^{63} were allowed to sponsor sacrifices^{64} — early forms of the inclusionism which later on characterizes Hinduism, in fact, until today.^{65}

The effect was the creation of a permament, and now, after all the liberal R̊gvedic intermingling and acculturation, of an artificial boundary between Aryans and non-Aryans (śūdra-ārya). The changes in the social formation, from semi-nomadic tribe to a larger tribal union, need a definite expression in order to be able to function as a “new order.” This is frequently expressed as antagonism between classes and groups both in language and ritual, and is first met with in the “first constitution of India,”^{66} the Purus.a hymn of R̊V 10.90. Such demarcations are, as can also be observed elsewhere,^{67} a typical reaction to an initial stage of free and wide-spread acculturation. The dominant brahma-ks.atra elite, already thoroughly mixed with local and aboriginal elements, now encapsuled itself vis-à-vis the “third estate”, the Vaiśyas, and stressed its superiority with regard to them, as well religious and racial “purity” over the non-Aryan Śūdras.

One of the strategies of the Kuru kings by which they achieved their new status was the traditional gaining of booty in their external expeditions (see below) and its distribution, but this is now supplemented by the collection of “taxes”, or rather, the coercion of “tribute”, bali. The Kuru king is the ideal type of a “benevolent lord” who seems to give more than he takes from his subjects and who supports his nobles and other subjects.^{68}

Sociologically speaking, this is typical for many early societies, whether based on collecting/hunting, simple horticulture or agriculture, or on (semi-)nomadic pastoralism. However, since the Kuru period, this kind of exchange, implemented throughout the realm, has been institutionalized in Indian society in a semi-religious fashion. As W. Rau has pointed out, the mutual relationship is expressed in the Vedic period by the concept of bhartr̊ :: bhārya “supporter/supported one” = attr̊ :: annam “eater/food”. The form of exchange follows a complicated pattern, a “social contract” that cannot be detailed here.^{69}

58See Kuiper, Rigvedic Loanwords, Studia Indologica, Fs. für W. Kirfel, Bonn 1955, p. 137-185, and Aryans in the R_{}˚ gveda, 1991; M.B. Emeneau, Language and Linguistic Area: Essays, Stanford 1980.

59For the term and concept, see M. N. Srinivas, Religion and Society among the Coorgs of South India. Oxford 1952; M. N. Srinivas, The Cohesive Role of Sanskritization, Delhi 1989, 56–72, J. F. Staal, Sanskrit and Sanskritization, Journal of Asian Studies, 22, 1963, 261 sqq.

60As Claessen calls it, Current Anthropology Aug.–Oct. 1984, 365–379; H.J.M. Claessen and P. Skalnik, The Early State. The Hague, 1978.

61See below, notes 93–95.

62Cf. the parallel in the development of the classs system in the Greek polis, see G. Nagy, Greek Mythology and Poetics, Ithaca, 1990.

63See Minkowski, The Rathakāra’s Eligibility to Sacrifice, IIJ 32, 1989, 177–194.

64See MS 2.9.5 on their chieftain, the nis.āda-sthapati, cf. KS 17.13, TS 4.5.4.2. AB 8.11 etc.

65Similar patterns of Hinduization are visible in modern Nepal, Orissa, in medieval Bali.

66See Paul Mus.

67The British in India did not react differently in the 1830s when their free-wheeling lifestyle as half-Indianized traders was changed by a series of “reforms” which brought them in line with the (soon to be “Victorian”) norms of their homeland.

68It has always belonged to the traditional duties of a king (and a ks.atriya) to collect and to distribute wealth, down to modern times: the king is supposed to distribute dāna.

The R̊VKh Kuntāpa hymns still reflect something of the old ideal in their description of the golden age of Pariks.it with the distribution of booty (vidatha) at a great festival about the time of the winter solstice. But the R̊gvedic pattern of a ritual exchange of goods and booty within a small tribe is now replaced by complicated (Śrauta) ritual and social exchange within the larger Kuru realm, in which, nevertheless, tribal sub-units survive. As has been pointed out, the Kurus had three, and their neighbors, the Pañcāla, six (originally three only). The great royal rituals underline the new and strengthened position of the king: Vedic ritual is not always as private as some think.^{70} The power of the Kuru king was qualitatively different from that and much greater than that of a chieftain, say of the Yadu tribe, in the R̊V (see below).

The expanded rituals are supported by the increasing stratification of society during the Mantra and YV Sam .hitā period. It is visible, apart from the establishment of the four classes (varn.a), in the formation of a large number of artisan specialists who are mentioned in the more complicated royal rituals such as the aśvamedha.^{71} This development coincides with an increasing production of goods: now, also the land between the rivers is settled and production increases;^{72} later on, the east is ‘reformed’ by the Kuru-Pañcāla Brahmins, whose ‘culture hero’, Agni Vaiśvānara, “sweetened the country as to make it suitable for agriculture”. Excess production apparently took place only after the establishment of a central power, such as that of the Kuru, that of the legendary Videha king, as well as that of the better attested Kosala and Magadha realms. Centralized power also brought about the perceived necessity of specialized crafts, best visible at the occasion of the great state rituals.

As far as the rest of the nobility and gentry was concerned, competition for superiority (śres.- t.hah. svānām, aham .śres.t.ha) among them was stimulated when it did not interfere with the role of the supreme chieftain, the king of the Kuru. All important positions in society were occupied by the alliance of Brahmins and Ks.atriyas (brahma-ks.atra), which exploited — according to their own words — the rest of the population; but within this new system competition was possible and indeed persistant.^{73} But rivalry and competition are also clearly visible in the newly developed Śrauta ritual.

Indeed, one of the strategies of achieving their goal of an internal competition without peril was the setting up by the Kuru kings of the complicated Śrauta ritual, which once and for all divided the people into four classes, and forged a new unity based on exchange between Brahmins and Ks.atriyas (brahma-ks.atra).^{74} The Kuru kings also succeeded in controlling, in turn, the older, amorphous groups of priests^{75} by a clear subdivision of their ritual labor. This was now re-distributed into four fields of specialization, i.e. the four Vedas and their ritual use. These four groups of priests (R̊V: hotr̊, SV: udgātr̊, YV: adhvaryu, AV: brahman)^{76} had quite specified duties during the solemn (Śrauta) rituals. The priestly occupations were even further divided, for good measure, into sub-specializations of 16 or 17 types^{77} — something not unlike the increasing specialization among the craftsmen and artisans. In both cases, centralized power stimulated specialization.

69Reminiscent of the modern jajmān relationships, but extending to other areas as well: it comprises a social contract including several generations, as well as the relationship between men, their ancestors and the gods; see a forthcoming publication, and in brief form, S. Jamison and M. Witzel in: A. Sharma, Hinduism (forthc.).

70Heesterman underlines their private character too much. Even a simple agnihotra can attract visitors — at least nowadays, and Vedic evidence points to similar, contemporaneous occurences. Certainly, larger rites, such as Soma sacrifice, and especially the horse sacrifice, were often disturbed by rivals.

71See e.g. the list in VādhB., = Caland, Kleine Schriften, Stuttgart 1990, p. 365

72See G. Erdosy, Urbanisation: The early settlements are on the banks of rivers, Yamunā and Gaṅgā, only; this is still reflected in some Mantras of the (much later) Gr_{}˚ hyasūtras, see Localisation p.205.

73The examples are too numerous to be quoted, see for example the many expression with bhrātr_{}˚ vya, and W.

Rau, Staat, p. 45.

With a political master stroke, the Kuru kings also succeeded in controling the aristocracy, that is their fellow Rājanyas and Ks.atriyas, by giving them something else, new, and fashionable to ‘worry about’: the complicated Śrauta ritual. In evaluating this one should disregard, for the moment, the usual phenomenological, pseudo-historical, and “philosophical” approaches to Vedic ritual and concentrate on its social effects.

It is surprising that even the contemporary specialists of Vedic ritual have not noticed that the Śrauta ritual — while often having one and the same aim, namely reaching heaven — is set up in such a way as to satisfy various levels of solemnity and stat us. A not (very wealthy) Vaiśya might have been content with the domestic (gr̊hya) rituals of passage that are executed for him and his family. However, a lower rank Ks.atriya might have attempted to go on to the next step on the socio-religious ladder and become a dīks.ita, that is an initiated “sacrificer” (yajamāna), after having learnt more of the Veda than a Vaiśya (such as a grāman.ī, a “trek leader”)^{78} or a lower rank Ks.atriya owner of such a wagon train (a grāmin), or a simple ks.atriyabandhu. After he had established the three sacred fires, he could then perform the Agnihotra, the New and Full Moon sacrifices, etc. If he wished for more, he could add the seasonal rituals (Cāturmāsya) and the yearly Soma ritual. If he was still not content with this and wished to impress his rivals further (who would often come to interfer with or destroy his rituals),^{79} he could go on with seven more types of soma rituals (soma-sam .sthā). While violent interference with one’s ritual may have been a remnant of a more agonistic period, as Heesterman believes, (this would be the one of the R̊V, not of a nebulous past!), ritual violence was still visible but tamed. Nobody takes the trouble to disturb a simple Agnihotra or New- and Full Moon ritual. It is the more important rituals, especially the Aśvamedha (capturing the horse), which bring out the rivals of the sponsor. What is important here is that these — only natural — rivalries were cleverly channeled in the new, Śrauta way of stratification.

74See W. Rau, Staat, p. 59–60, 118: they exploit the viś (an important item missing in R. Thapar, Lineages).

75See the enumerations in R_{}˚ V, with five or seven priests, e.g. R_{}˚ V 8.72: Adhvaryu, Hotr_{}˚ , etc.

76Even though the Atharvaveda remained a text which was not fully recognized for quite some time.

77Note the same technique in medieval Orissa where we find not one but four rājagurus, see G. Pfeffer, Puris Sasandörfer. Basis einer regionalen Elite (D. Litt. thesis), Heidelberg 1975. Pfeffer, Status and Affinity in Middle India, Wiesbaden 1982.

78See Rau, Staat p. 56.

79This was overly stressed by Heesterman; note that in even in the classical Vedic sattra ritual everyone could come and dispute with the sacrificers until the 11th day, see Falk, Bruderschaft und Würfelspiel, Freiburg 1986, p.

35.

Beyond the Ks.atriyas, the next level is that of the nobility of royal blood, i.e. of men who are “fit to be ordained as kings” (abhis.ecanīya^{80}), then that of the rājans themselves, not “kings” but rather “chieftains”; for example one of the 3 among the Kurus or one of the 6 of the Pañcāla. And, finally, there is the Great Chief, the King of the Kurus. The nobility had the means — and apparently also the leisure — to perform such rituals as the Agnicayana, a complicated rite taking a whole year, or — instead of the seasonal offerings (Cāturmāsya), — the gavām ayana which also takes a year. In similar gradation, a low rank ruler could receive, as pointed out above, the consecration as chieftain through the simple royal abhis.eka,^{81} the more complicated rājābhis.eka^{82} and ekarājābhis.eka,^{83} or the solemn aindrābhis.eka,^{84} and finally, there was the solemn Śrauta option of the rājasūya. Later on, a revised, complicated Śrauta version of the R̊gvedic, originally even Indo-European,^{85} horse sacrifice (aśvamedha), was added for especially powerful supreme kings who claimed “world domination,”^{86} which nevertheless only encompassed parts of (northern) India.

In the case of Pariks.it and his dynasty we can see the process of Śrauta adaptation happen before our eyes: It has not been noticed so far^{87} that another Mantra time text, Paippalāda Sam .- hitā 10, was composed to serve as “coronation” text of the early Kuru kings. It is here that we get for the first time the mentioning of typical Śrauta terms such as sava. More importantly, there is the connection, established by H. Falk, of the rājasūya and royal adoption, which is actually hinted at already in an older text, the Rohita book of AV (ŚS 13, PS 18.15-26). The son of the Aiks.vāku King Hariścandra also was called “the Red One” (Rohita, AB 7.13 sqq. in a story discussing the rājasūya); he apparently was added to this story in order to show his descendence from Rohita, the Sun (i.e. Vivasvant/Mārtān.d.a, one of the great Āditya gods; note the “solar” lineage of the Raghu/Rāma dynasty in the Rāmāyan.a; in fact, rohita was engendered with the help of another Āditya god, Varun.a). The term rohita also hints at the close connection of the “brilliance of the sun” and of royal glory (varcas, Avest. xvarenah). This whole complex, too, is in need of further investigation.^{88}

Summing up the discussion of ritual it can be said that by the time of the Mantra period, there were, on all levels of Indo-Aryan society, several ritual options available to each man if he wished to attain fame and glory, kīrti and (brahma)varcas. Every Rājanya, Ks.atriya and Vaiśya could perform such solemn rites on an offering ground near his home.^{89} The new Śrauta ritual thus put everyone in his proper station and at his proper place: in the life cycle and in society, both during the period of one’s Veda study and the roaming about as Vrātya as well as during one’s time as “settled” householder (gr̊hastha). There was opportunity for each and everyone to gain higher status by having the Brahmins perform more and more elaborate rituals — instead of simply raiding one’s neighbors.

80See W. Rau, Staat, p. 84 sqq.

81cf. Rau, Staat p. 89.

82KS 37.9, TB 2.7.15–17, BaudhŚS (as mr_{}˚ tyusava) 18.16–19, ĀpŚS 22.28, etc.

83As found in the Atharvavedic manuals and in Kauśikasūtra, see Author, The coronation rituals of Nepal, Heritage of the Kathmandu Valley, ed. by N. Gutschow and A. Michaels. St. Augustin 1987, 417–467.

84In AB 7. 13 sqq., see Author, On the localisation of Vedic texts and schools.

85For a brief summary see J. Puhvel, Comparative Mythology, Baltimore 1989, p. 269–276.

86As cakravartin, etc., see now S. Tambiah, The Buddhist Conception of Universal King and its Manifestations in South and South-East Asia. The first attested case is that of the great chieftain Sudās, R_{}˚ V 3.53.11–12, and the Aśvamedha of the Pūru chief Purukutsa, R_{}˚ V 4.42.8–9.

87But now see Tsuchiyama, Veda.no várcas, Indotetsugaku Bukkyogaku 5, 1990, 67–80.

88Note that “brilliance” and glory often are linked in Indo-Iranian thought (cf. Iran. x^{v}ar^{e}nah); see preceeding note.

89He had to ask the king for permission, however, to use land for this.

In order to carry out many of the religious and social reforms mentioned so far and as to achieve the general purpose of overlordship in northern India,^{90} the Kuru kings also initiated a collection of the major poetic and ritual texts, — certainly intended to show their care for traditional lore and knowledge. The “trick” was to preserve the old but to institute some, often minute changes as to serve the new ruler’s goals. In the case of traditional Aryan lore, the aim was not only to collect all (suitable) texts but also to re-arrange them in a fashion suitable for the new goals. The old ritual hymns and some poetry were assembled in the R̊gveda-Sam .hitā, the major ritual Mantras and early (now lost) explanatory prose texts in an Ur-Yajurveda-Sam .hitā, the melodies sung during the Soma sacrifice in an Ur-SV-Sam .hitā, and the healing and other charms as well as speculative hymns, though all reworked by Āṅgirasa Brahmins, in an Ur-AV-Sam .hitā.

What could have motivated the late R̊gvedic and early Mantra time poets and “copyright owners” and priests to make major changes in text transmission and ritual performance? The ‘extraction’ of the often secret r̊c and mantra texts from their authors’ and owners’ clans of poets and priests cannot have happened without a certain amount of pressure. Traditional owners of the “copyright” to a certain hymn were not likely to divulge the exact text or to voluntarily give up all their exclusive rights to the collection of texts composed and customarily also transmitted by their family or clan. Therefore, the carrot of “joint ownership” by the newly formed Brāhman.a class (R̊V 10.90) or, at least by those Brahmins learning just one Veda by heart, had to be offered as well.

Indeed, the tradition of individual and clan-wise origin of each hymn was preserved by a complicated system of arrangement of the R̊gvedic hymns in the “collection” (Sam .hitā), which also took into account the author whose name must be mentioned to this very day before reciting a hymn.^{91} Thus, the goal of having a new text collection fit for Śrauta ritual was achieved by preserving much of the traditional status of the poets/priests, their rights, and their “ownership” of compositions in sacred speech.

The collection of texts was not only made from the poets’ clans closely allied with the Bharata royal family, such as the Vāsis.t.ha, but in order not to lose continuity, also with hymns linked to the glorious past of the Pūru (and occasionally, even the Yadu-Turvaśa, Anu-Druhyu tribes). Many if not most of the traditionally remembered old hymns were included in the “national” collection of hymns, the R̊gveda, though the hymns of the Bharata and the Pūru clearly dominate the collection.

Once the collection was fixed, there was no need, of course, to create new hymns — which was a major goal of poets/priests (brahmán) in R̊gvedic times. What was still carried on was the composition of new speculative hymns: in the late R̊V, under the Bharatas, and especially in the AV under the Kurus. Note that the poet (brahmán) of the R̊V now reappeared as author of (part of the) AV, which was at first called Ātharva-Āṅgirasa, “the (collection of hymns) of the Atharvans and Aṅgiras.”^{92} Many of these new hymns deal with the ritual and its “philosophical” underpinnings, such as yajña, “food” (brahma-)odana, ucchis.t.a, etc. which are in need of a detailed investigation.

90See ŚB 13.5.4.1 sqq. with hymns which praise the Bharatas, especially 13.5.4.23.

91The arrangement is based on the author, deity, type of meter and length of hymns and the number of hymns in each collection that originally belonged to one poets’ clan.

In all these cases one can notice that one means to bring about continuity in spite of the great changes carried out under the Kurus, was the artificial archaization of certain parts of the new Śrauta ritual,^{93} the use of artificial, archaic forms^{94} in the poetic and learned language of the poets, priests and “theologians” of the Mantra and YV Sam .hitā periods, and of text formation and their collection.^{95} The new ritual and its language appeared to be more elaborate and impressive but at the same time, had to give the appearance of having come down from a hallowed past.

The formation of early states such as that of the Kurus usually brings about important changes in ideology, religion and mythology. The new religious and political ideology necessary for the expanded dimension in tribal organization included many elements of the older, R̊gvedic beliefs about mankind’s descent from the gods, their access to heaven and to eternal “happiness” after death: for example, warriors who died in battle were taken to heaven by the Apsaras (the rather unnoticed sisters of the Norse Walkyries), and they were readily accepted there by the gods — a topic found from the R̊gveda^{96} onwards.^{97} Access to heaven is one of the major topics of all Śrauta ritual. Next to sons, rain, cattle, long life (āyus) it is a prolonged (theoretically, but only by exception, eternal) stay in heaven that one strives for, after a stint on earth characterized by constant strife and frequent hunger, as has been described in detail by W. Rau.^{98} It is in the Kuruks.etra area that the heavenly river (Sarasvatī, the Milky Way) was regarded as flowing down from heaven about the time of the winter solstice: it thus opened, with its two branches touching

92Most of the authors of the R_{}˚ gvedic hymns belong to the Aṅgiras clans.

93For details see my forthcoming monograph: The Realm of the Kurus. Only the results of this investigation can be summarized here. As for the ritual, note that the metal (ayasmaya) vessel of the R_{}˚ gvedic pravargya (or gharma, mahāvīra) rite, R_{}˚ V 5.30.12–14 etc., was substituted by a clay one fashioned in a very primitive, archaic way by the Brahmin priests themselves; ritual garments were freshly made in a very archaic fashion, without the use of needle and thread; or, in the reorganization of the labor of the priests many old names of priests and their job descriptions were kept, while the role of the Brahman changes markedly.

94R_{}˚ gvedic kr_{}˚n .oti, kr_{}˚n .umah., kr_{}˚n .u (PS) for the atharvavedic (and later) allegro forms karoti, kurmah., kuru, see K. Hoffmann, AzI p. 575 sqq.; — PS kr_{}˚n .va is more archaic than the R_{}˚ gvedic form kan .va, see K. Hoffmann, AzI, 15 ff.: writing in 1940, he could not yet know but reconstructed the correct form kr_{}˚n .va (PS) as predecessor of kan .va. — Archaic forms are also attested to some extent in another Kuru text, KS (and much less so in TS). — The artificial Taitt. form suvar may be an imitation of older Kuru archaisms. Note that we find suvar, suvarga but svasti; further: the rather artificial introduction of this principle in uv eva for u eva/-v eva: See Tracing p. 174, 178. — Note also the l/r treatment; R_{}˚ V 10 has the popular l where KS, PS often avoid them, e.g. rohita, romaśa, ariklava, see Author, Notes on Vedic Dialects, 1, Zimbun 67, p. 44 sq.

95Inclusion of all the r_{}˚ c materials, including some old Yadu-Turvaśa, Anu-Druhyu and many of the Pūru hymns; note also the archaization of Atharvaveda Mantras by starting hymns with hieratic meters. — Oldenberg (Prolegomena) has noticed that the archaizing tendency further increases with time: the later SV, YV texts tend to introduce R_{}˚ gvedic forms instead of their own traditional ones.

96Note the promise given to Purūravas, R_{}˚ V 10.90.18; cf. KU 1.4, etc.

97All through classical literature, even as late as 1149 A.D. in the Rājataraṅgin.ī where the motif figures prominently: Rāj. 7.173 “brave men wedded to the heavenly maids (Apsaras)...” The concept is perpetuated in the later Rājataraṅgin.īs of Jonarāja, etc. — Cf. also Bhagavadgītā 2.31–32 (yudhāc chreyo ’nyat ks.atriyasya na vidyate // yadr_{}˚ cchayā copapannam . svargadvāram apavr_{}˚ tam / sukhinah. ks.atriyāh. ... labhante yuddham īdr_{}˚śam).

98Rau, Staat, p. 31.

the north-eastern horizon, the “gate” to heaven.^{99}

An important, if not the chief one among the religious developments is that the new royal center in Kuruks.etra gave rise to a new mythology of the region. It is here and not elsewhere that the gods traditionally sacrifice and hold their long sattra rites to overcome their perpetual foes, the asuras.^{100} Further, the river Sarasvatī itself is the personification on Earth of the goddess Sarasvatī, the name of the Milky Way in the Vedic texts; this falls down on earth at the Plaks.a Prāsravan.a, the world tree at the center of heaven and earth,^{101} and then continues to flow through the land of the Kuru people,^{102} — which is identified with the whole earth.^{103} The area was conceived as the “center of the world”, a trait first visible after the victory of the Bharata king Sudās and his settling on the Sarasvatī (located according to R̊V 3.53.11 vara ā pr̊thivyāh. “at the best place of the earth.”^{104}) Such identifications of one’s habitat with the center of the world are common among many peoples, and in such cases the place of the ritual always is regarded as the center;^{105} However, Kuruks.etra now also became the place where even the gods offer (devayajana).^{106}

99This is a transformation of the older R_{}˚ gvedic concept of the heavenly gates, a standard feature found in the Āprī hymns, see L. van den Bosch, The Āprī hymns of the R_{}˚ gveda and their interpretation, IIJ 28, 1985, 95–122, 169–122, and cf. Author, Sur le chemin du ciel. Bulletin des Etudes indiennes, 2, 1984, 213–279. — The old idea of an inverted tree is found in the R_{}˚ gvedic aśvattha, held upside down by Varun.a 1.24.7, cf. TĀ 1.11.5, Kat.hUp. 6.1, BhG 15.1-3 (see Kuiper, The Heavenly Bucket, in: India Maior (Fs. Gonda), Leiden 1972, 144–156, cf. The Bliss of aša, IIJ 8, 1964, 117; and see now Author, Looking for the Heavenly Cascet, EJVS 1-2 and a more detailed version in: Fs. Thieme, StII 1996, forthc.). By performing a sort of pilgrimage along its banks, against the current of the stream, one could move along the Sarasvatī (= Milky Way) through the night sky and reach heaven. 100PB 25.13 “Indra and Ruśamā made a wager: ‘Whichever of us shall first run around the earth shall be the winner.’ Indra ran around the earth, Ruśamā ran around Kuruks.etra (only).” This indicates that Kuruks.etra is identified with the whole earth; cf. Hoffmann, Aufsätze zur Indo-Iranistik, ed. J. Narten, Vol. 1, Wiesbaden 1975, p.7. 101See Author, Sur le chemin du ciel; — one span north of Plaks.a Prāsravan.a (the Plaks.a tree of “forth-streaming”) is the center of both heaven and earth (JB, VādhPiS); cf. in Germanic lore, Yggdrasil, which is the source of three rivers. 102The concept of Kuruks.etra occurs first at MS 2.1.4, 4.5.9; then at AB 7.30, ŚB, JB, PB, Kat.hB, and TĀ 5.1.1 (with a description of its borders); that of devayajana as early as ŚS 10.5.15-20, PS 16.129.1-5, MS 1.1.8, KS 1.7, TS 1.2.3., etc. but the two are not immediately linked formally; though descriptions such as MS 2.1.4:5.9 are clear enough: dev´ā vái sattrám āsata kuruks.etré. 103Later ritual (PB 25.13, JB, LŚS, etc.) makes the Sarasvatī and her companion river, the Dr_{}˚śadvatī, the place of long treks which one may perhaps call “pilgrimages” along their banks, leading to heaven, as the heavenly river (Sarasvatī, the Milky Way) was regarded as flowing down from heaven at this spot about the time of the winter solstice: it thus opened, with its two branches touching the north-eastern horizon, the “gate” to heaven. — All of this is unmatched by any other area mentioned in Vedic texts; places like Prayāga and Kāśi or even the Naimis.a forest (though mentioned already in KS 10.6:130.8) attain this kind of fame only well after the Vedic period. 104“At the navel of the earth” R_{}˚ V 8.43.4 nabhā pr_{}˚ thivyā, bhuvanasya majmane; and “the best place on earth” R_{}˚ V 3.23.4 ní tvā dadhe vára ´ ā pr_{}˚ thivy´ā, íl.āyās padé; 3.53.11 Sudās will offer at the best place on earth: athā yajāte vara ā pr_{}˚ thivyāh.. 105At this location, the Plaks.a tree clearly is the central world tree which pushes up heaven. See Author, Eine fünfte Mitteilung über das Vādhūla-Sūtra. StII 1, 75–108, and Sur le chemin du ciel. Bulletin des Etudes indiennes, 2, 213–279; 106MS 2.1.4:5.9 dev´ā vái sattrám āsata kuruks.etré.

IV. STRUCTURE

The immediate outcome of the establishment of the new system of Śrauta ritual for a king of the Kurus was: his “reform” unified various smaller tribes by a single, but complicated network of mutual ritual relations; this frequently was of a dualistic and partly antagonistic nature (note especially the Kuru-Pañcāla vrātya relationship which imitates the deva :: asura strife in myth, and the ārya :: śūdra competition in society).^{107}

The older dual organization of the Five Peoples of R̊gvedic times (Anu-Druhyu, Yadu-Tur-vaśa) was probably echoed in the Kuru Kingdom, originally, by that of the Bharata-Pūru. This pattern emerges more clearly when the Kurus started to spread eastwards. The new territories (up to Kausambi / Allahabad) were settled by groups who then organized themselves as the Pañcāla tribe^{108} which was explicitly divided into six sub-units, — a fact which should lay to rest all speculation about the origin of the name in the number “five”.^{109} The (Pūru-)Iks.vāku are mentioned to have settled on their eastern border near Benares.^{110}

The relationship between Kurus and Pañcālas was ambiguous. On the one hand, both royal families intermarried.^{111} This, actually, was one of the strategies of the supreme Kuru king aiming at asserting his authority at this highest level, and has been a favorite method in all early states which cannot rely on paid bureaucrats but must rely on various types of relations built on personal loyalty between the ruler and his nobles. Polygamy, which is well attested for Vedic kings,^{112} helped to establish multiple relationships with important external and internal noble families, something which certainly was necessary as the Kurus and Pañcālas still were divided into three viz. six powerful subgroups. The device, in fact, was one of the means to forge alliances between various exogamic units of gotras even for the richer ones among the Ks.atriyas and Brahmins.^{113}

On the other hand, the union (of ritual^{114} moieties) of the Kurus and Pañcālas was stressed by the custom of sending their Männerbund associations (vrātya) into each other’s territory: note the KS 10.6 story about king Dhr̊tarās.t.ra Vaicitravīrya,^{115} and also the story in BŚS 18.26 about the vrātyas of the Kurus in the land of the Pañcālas.^{116} Most telling, perhaps, is the note about the Southern (Madhya Pradesh) people at JB 3.146 who send their sons northwards to the Kurus.^{117} It is important to note that these young men associations do not enter new, “virgin” territory south of the Vindhya but go, in their vrātya excursions, just as the Kurus and Pañcālas respectively, to the land of their closest orthoprax neighbors. This clearly indicates that vrātya exchange is carried out between (nominal, ritual) allies,^{118} and not between, for example, the Aryans and the aboriginal tribes of the south. — At the same time, these stories indicate something of the traditional aggression resulting in cattle rustling, fighting and small scale warfare existing with one’s neighbors which was now canalized by the new (vrātya) ritual.

107Note that it is formalized in the mahāvrata ritual as a śūdra :: ārya conflict (ĀpŚS 21.17 sqq. etc.); the discussion on the aśvamedha mentions the taking away the goods of the lower classes, see Vādhūla Śrautasūtra 3.79 = Kl. Schr. p. 370 sq. — During the R_{}˚ gvedic period it might have been a Bharata :: Pūru configuration.

108This included the R_{}˚ gvedic Krivi tribe, see above for other members.

109See literature in Mayrhofer, KEWA II 188, EWA II, 66. — Note that the YV school of the Pañcāla, the Taittirīya, has 6 subschools (Baudh., Vādh., Bhār., Āp., Hir., Vaikh.) which echo, also in location (see Author, Localisation, p. 205), the division of this great tribe.

110Iks.vāku in the eastern parts of the Pañcāla-Kosala area, see Tracing, n. 253, 349, 389; cf. JB 3.168–170: par.

190. Note that JB has Tryarun.a as an Iks.vāku, see Author, Rigvedic history (above, n. 15), par. 5c with n. 57; cf.

also JB 3.237–238: par.204, JUB 4.6.1.2.

111See JB 2.278–9.

112Note the technical names of the Kuru king’s wives already in the Kuntāpa hymns, R_{}˚ VKh.

113A well known example are the two wives, Maitreyī and Kātyāyanī, of the rich (mahāśālina) Yājñavalkya.

Polygamy led to the necessity to specify the mother of a certain prince (or a Brahmin’s son): See the names in -putra in the Maurya, Sātavahana dynasties, and of the authors of ŚB in the Vam .śas of this text.

114Note the ritual competition between the two tribes, e.g. at JB 1.262 where they hold a debate (brahmodya) on cosmological and theological questions.

115The Vrātyas were not accepted by him, and therefore destroyed his cattle with the help of some rituals. Why this Kuru king? Does this reflect the Vrātya, i.e. a not always amiable relation between Kaurava and the Kuru-Pañcāla Vrātya (here led by a Pañcāla: Vaka Dālbhi). See Falk, Bruderschaft, p. 58 sqq.

Through the ritualization of these relationships, all social and political energy could now be projected either towards a common goal, namely expansion to the east and south,^{119} or otherwise towards more or less innocuous, often petty rituals which enabled ambitious Ks.atriyas to compete with their neighbors and rivals. This sort of rivalry always existed, even within the clan, where one wanted to become śres.t.hah. svānām.

While the stategies described so far were successful in the establishment and enlargement of Kuru power, the rather minute origins of the Kuru realm are still reflected by the terms used for its incipient “administration.” The titles of the royal functionaries^{120} are designations such as ‘butcher’, or ‘dice-thrower’, ‘meat cutter’, etc. At first glance, these seem to be rather minor servants at the home of the king.

The originally quite small nucleus of Kuru power is also visible in the (royal) ritual itself. Most of the older, R̊gvedic and tribal rituals were linked to the course of the days and nights, the phases of the moon, the seasons, and the course of the sun. They are “rites of passage” of the year. However, the new unified and rearranged Śrauta ritual, with its highly archaizing tendency,^{121} not only included all aspects and all officiants of the older rituals, but it also included some major new royal and ‘national’ rituals. These took place, just as their counterparts performed by the gods, at the Kurus’ spiritual and political center in Kuruks.etra, at Āsandivat ‘the (place) having the throne’, obviously a (temporary) seat of the frequently travelling Kuru kings;^{122} other names for their royal settlement (sādana) are Nad.apit and Rohītakūla.^{123} Some other rites took place “at the back of Kuruks.etra” at Parisaraka or Parīnah^{124} where the river Sarasvatī disappeared in the desert.

116See Falk, Bruderschaft, p. 55 sqq.

117Cf. Rau, Staat p. 14 “wenn ein Vater seinen Sohn aussiedelt, dann siedelt er ihn im Norden aus,” which Rau understands as settlement. I think this rather is a question of Vrātya movement to one’s neighbors.

118The “south”, i.e. the JB territory north of the Vindhya and south of the Yamunā, is inhabited by Matsya (on the Yamunā), by the Satvants, both going back to R_{}˚ gvedic times, and apparently also by the (aboriginal?) Kunti, cf. MS 4.2.6 Kúrūn .ām . kaunté. — Note the fight of the Pañcālas with the Kuntis, see Author, Tracing, n. 113, KS 26.9, end.

119See KS 26.2:123.17, MS 4.7.9:104.14, TB 1.8.4.1, but contrast ŚB 5.5.2.3–5, cf. and W. Rau, Staat p. 13, Author, Localisation p. 178, Bodewitz, transl. JB 1.66 sqq., p.276 n. 31. — Note that this also refers to ritual: JB 1.262:94 is very clear in this regard: the Śrauta ritual had by then spread even to the udantyas, see Localisation p.

187.

120See W. Rau, Staat, 107 sqq.; now H. Scharfe, The State in Indian Tradition, Leiden 1989; the officials (ratna/ratnin) are: purohita: house chaplain, senānī: army leader, grāman .ī: smaller leader or later, village head, sūta: herald, sam .grahītr_{}˚ : chariot driver, ks.attr_{}˚ : meat cutter, bhāgadugha: food distributor at the ‘table’, govyaccha / govikarta:

butcher, aks.avāpa: dice thrower, taks.an: carpenter, rathakāra: chariot maker, pālāgala: runner, and other rājopajīvins; also the sthapati of the aboriginees (Rau, Staat p.113): of Nis.āda, others; ugra/pratyenas “henchmen”

(who do not belong to aristocracy); cf. also: dūta “emissary”, abhikrośa “herald”, piśuna, pariskanda“?”

121See above, n. 93; cf. below ad n. 165, 167.

122Note that the kings roam about in their territory because of their comparative lack of centralized power, in

All of these strategies and the changes brought about underline the increased power of the supreme rājan of the Kuru as a new “great chief”. The relation between the royal court, the subtribes (jana, janatā), the clans (gotra), and the individual families (kula) was characterized by the ability of the higher levels in the social hierarchy to extract tribute (bali) from the lower levels. These tributes (in kind) still were to some degree “recycled” during the great rituals just as they had been in R̊gvedic times (during the vidatha^{125}). However, the royal officials of the budding administration of the Kuru kings also took their “fair” share. That this was not always acquired in genial fashion can be noticed already in a Mantra time text, at Atharvaveda 3.29.1, which describes the other world as one where one has to give up just one sixteenth^{126} as tribute. The Mantra and Brāhman.a texts bear frequent witness to the relatively undetermined and arbitrary nature of this kind of “tax”.^{127} Sahlins has described this type of society in some detail.^{128}

The royal officals were “paid” by the king from his bali. They did not hold just ceremonial offices (such as the govikartā, etc.^{129}) but had real administrative functions as well: as army leader (senānī), herold (kāru) or emisssary (dūta, sūta), and as royal priest (purohita) who was closely linked to the actual carrying out of government and who was very closely allied to the King also on a personal level, sometimes as his chariot driver.

The various levels of authority within the new Kuru super-tribe are discernable to some degree: At the top was the king (ekarājan^{130} as a JB 2.275 and later an Atharvavedic text have it), his relatives and his peers (the high aristocray), from which a lone the king could be chosen (rājya).^{131} Below this ranged the smaller chiefs (rājan, three in the Kuru, and six in the Pañcāla tribe).^{132} Then came the leaders of the various clans who strove to become “the best” (bhrātr̊vya, aham .- śres.t.ha); for them a title is not found. They may, however, often have been identical with the owner (grāmin) viz. the leader of a wagon train^{133} (grāman.ī). Significantly, this term was first introduced in R̊V 10, thus under the Bharata chieftains. Finally, there is the head of the extended family (dam .pati, pitā(maha), pati).^{134}

order to control the various parts of their realm, cf. Claessen, Current Anthropology Aug.–Oct. 1984, 365–379, Rau, Staat, p. 128.

123PB 14.3.11–12; cf. JB 3.183: par. 192: the high bank (kūla) of the Yamunā; cf. also rohita PS 18.15-26 = AV 13, and his relation to the (red) sun, varcas (cf. x^{v}ar^{e}nah).

124See Author, Sur le chemin du ciel.

125See Kuiper, The ancient Aryan verbal contest, IIJ 4, 1960, pp. 217–281.

126See already R_{}˚ V 8.47.17; AV 3.29.1 “What the kings share among themselves — the sixteenth of what has been offered-and-bestowed...” (Whitney).

127Examples in W. Rau, Staat, par. 24.3, 25, 40.2, p. 104.

128Sahlins, Poor Man, Rich Man, Big Man, Chief: Political Types in Melanesia and Polynesia, Comparative Studies in Society and History 5, 1963, p. 285–303.

129Rau, Staat, p. 107 130In the late Kauśikasūtra; cf. the Mantra time names samrāj, adhirāj see Author, Localisation p. 183.

131See Rau, Staat, p. 68: The Vaiśya and others were a-rājya.

132Note that according to Rau, Staat, one could be a rājan even before one’s “coronation”. Note also the more than 7000 “rājas” of the Vesali people in the Pāli texts. — Cf. S. Zimmer, viś´ām páti und viśpati, MSS 44, 291–304.

133Cf. the Germanic term Germ. Herzog, Dutch hartog “the one who leads in front of the army [on the move]”.

See now H. Kulke on the wish for a settlement: Grāmakāma — “das Verlangen nach einem Dorf”. Überlegungen zum Beginn frühstaatlicher Entwicklung im vedischen Indien. Saeculum Bd. 42, 1991, 111–128.

The king could exert his will by a ready band of “terrible [warriors]” (ugra)^{135} or henchmen. He also relied on a network of spies, known since R̊gvedic times as spaś, in the Brāhman.as perhaps as piśuna; this institution was perfected under the early empires, as described in detail by Kaut.alya (as cāra). Nevertheless, the chieftain and even the great chief of the Kuru, was not yet, by any means, an absolute monarch. He could be disposed by a rebellion among his peers or by the people. This happened fairly frequently; the person of the exiled king is a recurrent in the texts of the YV Sam .hitā and the Brāhman.as, and special rituals were created to let him regain his kingdom.^{136} The new concentration of power created, as Sahlins notes,^{137} at the same time, the roots for its destruction, of revolution. Absolute power was realized only in the first great states with aspirations of empire, such as Magadha about 500 B.C.^{138} The Vedic Kuru realm still resembles that of a large Polynesian chieftainship^{139} such as that of Hawaii — and with a similar ideology.^{140} In its origin and size, though not in its ideology,^{141} it may be compared with another early state, the realm of the Franks under the Merowingian kings.

The new arrangement of the Vedic society — superficially united in a diversity of four classes — did not only provide each member of the new Kuru super-tribe with a clear and fixed identity but it also allowed society to eliminate much of intra-tribal and inter-tribal strife, such as the constant cattle rustling, and to turn the Ks.atriyas’ activities outwards. Military expansion quickly established the new Kuru tribe as the only major force among the few remaining smaller tribes of Northern India, such as the Matsya, Satvant, Uśīnara. While the nuclear area of the Kuru was the eastern Panjab, Haryana and the western part of Uttar Pradesh, the Kurus soon made their presence felt beyond this.

The military expansion of the Kurus may have been limited to the periodic raiding and looting of new agricultural crops,^{142} taken from the pockets of aboriginal agricultural populations (Nis.āda, “those sitting at their proper places”), and may have occurred as recurrent parts of the yearly transhumance movements. But in other cases^{143} we notice a long-range advance. The texts clearly describe this as happening in two directions: eastwards (KS 26.2:123.17; including the victory over the Kāśi^{144}) and southwards (MS 4.7.9:104.14).^{145} This means, expansion into the territories of the materially little progressed, chalcolithic cultures of the east and into the lands of the aboriginal agricultural peoples of the south, the Banas/Malwa cultures along the Chambal river.

134See Rau, Staat, p. 38 sqq. for details of the later, YV Sam .hitā and the Brāhman.a period.

135See W. Rau, Staat p. 114, who takes them as some sort of military men, refering to the famous passage of BĀU 3.8.2; they were ranked higher than other officials such as the grāman .ī, etc. Cf. also pratyenas, loc. cit.

136See W. Rau, Staat, on revolts and the exiled king, p. 128 sq., e.g. Balhika Prātipīya, the Kauravya king (ŚB 12.9.3.1 sqq.) regains the chieftainship over the Kuru subtribe of the Sr_{}˚ñjaya.

137Sahlins, Poor Man, Rich Man, Big Man, Chief (see above, n. 128).

138Which is, consequently, characterized in quite unpleasant fashion by the Indian Epic in the person of king Jārasandha of Magadha (MBh. 2.14 sqq.).

139Sahlins has described this type of society in some detail for Polynesia.

140Which is not only an accident of history or the result of social and economic development. Both societies make use of the old Eurasian “ideology” of a descent of the chiefs from the Sun deity. More on this see Author, Vala and Iwato. The myth of the hidden sun, (forthcoming).

141Though the Vedic (and later Indian) kings had to give (dāna) they did not and could not hand out such gifts continually as the Franconian kings did (cf. Claessen, Current Anthropoplogy 1984, 365 sqq). Rather, they had to rely on periodically making booty in the east and of distributing new pasture (and also fields for barley and rice agriculture) to their nobility and their people (viś).

142See W. Rau, Staat, p.13, and Heesterman, Contributions to Indian Sociology, N.S. 15, 258, ad TB 1.8.4.1.

143Such as the march eastwards of the Kurus (TB 1.8.4.1), the victory of the Pañcālas over the Kāśi, ŚB 13.5.4.19.

As confirmed by the development of the Vedic dialects,^{146} the newly stratified society of the Kurus with its model of orthopraxy emerged from a fairly small territorial nucleus and the new pattern spread quickly in all directions, as far as the natural conditions of the subcontinent would allow. The testimony of the texts, their language, and archaeology indicate the expansion of these traits from a small nucleus centering around Kuruks.etra.

Considerable re-organization of texts, rituals and social functions therefore were the hallmark of the strategies underlying the establishment of the Kuru realm. Especially the introduction of the Śrauta type ritual, the division of labor between the King and his brahmins, the close cooperation between Ks.atriyas and Brahmins (brahmaks.atra), and even more so, the establishment of the system of four “classes” (varn.a) was to become seminal for the development of Indian society ever since.

***

The Kuru realm survived under Pariks.it’s descendant, Janamejaya Pāriks.ita, Janamejaya’s sons, his grandson Augrasainya, and probably beyond this.^{147} A closer reading of the texts yields more results for this still very hazy picture of Vedic history: the Kurus were overcome by the (probably non-Vedic) Salvas^{148} who “dispersed the Kurus from Kuruks.etra” (JB 2.206, ŚŚS 15.16.11–12,^{149}) — a fact completely overlooked by the historians of old India. The Salva (or Salvi), mentioned at ĀpMp 2.11.12 as Yaugandhara, settle opposite or near the Matsyas on the Yamunā. By that time, the former and by now defeated Kuru tribe and the Salvas had apparently coalesced and they therefore re-appear in a late Brāhman.a^{150} as a standard Vedic people, and then, in the Epic and Buddhist literature also as Śūrasena.^{151} The name of the Kurus was kept alive during the Vedic period, and, in fact, the area appears to have been thoroughly (re-)Sanskritized^{152} already by the time of ŚB: the Kurus are not reckoned with the despised Bāhīkas (“the Outsiders”) of the Panjab but are again regarded as belonging to the heartland of orthopraxy.^{153}

144See ŚB 13.5.4.19 sqq.; note that Ajātaśatru is (at the same time?) king of the Kuru (VādhB) and of the Kāśi.

145Cf. Rau, Staat, p. 13, without making use of the different direction of movement, though. Author, Das Kat.ha Āran .yaka, Erlangen 1974 and Kulke, A History of India, 1986, make this connection.

146See Author, Tracing the Vedic Dialects.

147Cf. ŚB on the fights with the Kāśi.

148Later known as the Y(a)ugandhara: ĀpMp. 2.11.12.

149The fight of the Kurus with the Salvas may be echoed by the great battle of the Mahābhārata where it is confused with the R_{}˚ gvedic dāśarājña (between the Bharatas and the other tribes of the Panjab; note also the earlier Twenty Kings’ Battle of R_{}˚ V 1.5 3.9); all of this was distant memory even by the time of R_{}˚ V Khilas, and definitely so in JB 3.245: par. 205, KS 21.10:50.1, MS 3.7.7: 40.6.

150Cf. also ŚB 10.4.1.10, where the Salvas have the regular classes (varn .a): rājānah., brāhman .āh., and vaiśyāh..

151Perhaps originally, a nickname, cf. the names of the sons of Pariks.it, such as Ugrasena etc. — The name is foreshadowed in JB 1.262: “Therefore, among the Kuru-Pañcālas, a hero (vīra) is born with all the heroes.” — In the Epic they occasionally occur next to the Kuru(-Pañcālas).

152For the nature of the Salvas, see Author, Tracing, n.333; cf. further, the information of JB 2. 297–299 on Vyādhinīh., the female hunters which may be compared to the story of Alexander meeting the queen of the Amazones in north-west India, see now E. Garzilli, First Greek and Latin documents on sahagamana and some connected problems, with notes 105–106, IIJ

Against this background it is not surprising that a late Vedic text, BĀU 3.3.1, can look back at the royal family of the Kuru as belonging to the distant past. The passage (ŚB 14.6.3 = BĀUK 3.3.1-2) mentions the fate of the Pāriks.itas, the royal family of the Kurus, and asks: “what has become of the Pāriks.itas?” — The answer is: “they have, in truth, gone whither the offerers of horse sacrifice go” — and this is a ‘heaven’, a place beyond the ends of the world, and the ring ocean around it. Apparently they have gained a firm place in heaven which is otherwise granted only to such extra-ordinary persons such as the Seven R̊s.is.

These few passages might, ultimately, provide the clue for the prominence, in the later YV-Sam .hitā and the Brāhman.a period, of the Pañcālas with their Taittirīya, Kaus.ītaki, Śāt.yāyana schools. The political and cultural center now had shifted from the Kuru to this tribe which lived more towards the East, in what is now Uttar Pradesh. The Pañcāla king Keśin Dālbhya and his successors are prominent in a later YV-Sam .hitā, TS, and beyond.^{154}

According to JB 2.278-9, however, Keśin was closely related to the royal family of the Kuru: his maternal uncle was Ucchaiśravas, son of Kuvaya, the King of Kurus (kauravya rājā). Apparently he simply took over when the Kuru line was in decline (or without heirs?), due to the Salva invasion. Keśin is also credited with the ‘invention’ of the dīks.ā of the Soma sacrifice.^{155} He is both the new political as well as “spiritual” leader. The power of this dynasty lasted much longer than that of the original Pāriks.ita dynasty. His descendants are reported as being numerous even in the comparatively late ŚB where they are called Dālbhya/Dārbhya Kaiśina.^{156}

The mysterious Iks.vākus, which already appear at R̊V 10.60.4 may help to explain the developments in the Eastern part of the Pañcāla area. They are mentioned already in AV 19.39.9 = PS 7.10.9 as one of the Eastern groups (with the Kāśi and Matsya) living at the edge of Indo-Aryan settlements.^{157} In the Pāli texts (DN 3.1.15 sqq.), Okkāka (Skt. Iks.vāku) is the forefather of the Śākyas, who lived in the Central Tarai of Southern Nepal. A connection of the Iks.vāku territory with that of the Kān.vas (of the Brāhman.a period and their ŚBK texts) is highly probable. The Rāmāyan.a, of course, takes the Iks.vāku ancestry of the Kosala dynasty of Rāma for granted. Its appearence in the East may be directly correlated with the movement of the King Videgha Māthava into the country East of the Sadānīrā. (To the Kān.va, this is the country East of the Kuru-Pāñcālas, i.e. Kosala).^{158}

153Note also the fierce Kathaíoi “tribe” (i.e. Kat.ha Brahmins) who live in the same area as the Salva (and Mahāvr_{}˚s.a) at the time of Alexander, see Arrian, Anabasis 5.22. The “tribe” is reported to be one of the fiercest in north-west India. — The bliss of the Kuru realm is even remembered in Jātaka no. 276 of the Pāli canon, the Kurudhamma-Jātaka.

154For the prominence of Keśin Dārbhya, see Author, Notes on Vedic Dialects I, Zimbun 25, 1990, 40–41; cf. also JB 1.285 and cf. P.I. Koskikallio in the preceding issue of EJVS.

155Keśin’s invention of the consecration to the soma ritual, the Kaiśinī dīks.ā, is told in VādhB 4.37 = Caland, Kl.

Schr. 147 sq., KB 7.4, JB 2.53; cf. Sri Krishna Sharma, Keśin Dārbhya and the legend of his dīks.ā, ABORI 48/49, 1968, 241–245.

156ŚB 11.8.4.6 says that Keśin’s descendants continue to survive.

157Cf. also Iks.vāku in the Vādhūla Mantras, see Author, Early Eastern Iran and the Atharvaveda. Persica 9 (1980), pp. 86–128. Cf. on Iks.vāku myths, Berger, WZKSO 3, 1959, 34–82.

158Note the intention of the story: Gotama Rāhūgan.a is otherwise known only as the author of R_{}˚ gvedic hymns. To

The famous Videgha Māthava legend of ŚB 1.4.1.10 sqq. tells the story of the “civilization process of the East” in terms of its Brahmanical authors, and not, as usally termed, as the tale of “the Aryan move eastwards.^{159} For it is not only Videgha Māthava, a king living on the Sarasvatī, but also his priest Gotama Rāhūgan.a who move towards the east. Not only is the starting point of this “expedition” the holy land of Kuruks.etra; the royal priest, Gotama Rāhūgan.a, is a well known poet of R̊gvedic poems as well, — and thus, completely anachronistic.^{160} Further, the story expressively mentions the role of Agni Vaiśvānara, the ritual fire, in making the marshy country of the East arable and acceptable for Brahmins. All of this points to Sanskritization or rather, Brahmanization) and Ks.atriyazation^{161} rather than to military expansion.^{162}

The Māthavas, about whom nothing is known outside the ŚB, may be identical with the máthai of Megasthenes (c. 300 B.C.), who places them East of the Pazálai (Pañcāla), at the confluence of the Erénnesis (Son) with the Ganges.^{163} The movement of some clans, with their king Videgha and his Purohita, eastwards from the River Sarasvatī in Kuruks.etra towards Bihar thus represents the ‘ritual occupation’ of Kosala(-Videha) by the bearers of orthoprax (and orthodox) Kuru culture, but it does not represent an account of the first settlement of the East by Indo-Aryan speaking tribes which must have taken place much earlier as the (still scanty) materials of archaeology indeed indicate.^{164}

V. SUMMARY

It can be said that the Bhārata/Kaurava/Pāriks.ita dynasty of the Kurus sucessfully carried out and institutionalized a large scale re-organization of the old R̊gvedic society. Many aspects of the new ritual, of the learned speech, of the texts and their formation reflect the wish of the royal Kuru lineage and their Brahmins to be more archaic^{165} than much of the texts and rites they inherited. In this fashion, the new Pāriks.ita kings of the Kurus betray themselves as typical newcomers and upstarts who wanted to enhance their position in society through the well-known process of “Sanskritization.” In fact, to use this modern term out of its usual context, the establishment of the Kuru realm was accompanied by the First Sanskritization.^{166} Incipient state formation can only be aided if it is not combined with the overthrow of all inherited institutions, rituals, customs, and beliefs. The process is much more successful if one rather tries to bend them to one’s goals or tries to introduce smaller or larger modifications resulting in a totally new set-up. The new orthopraxy (and its accompanying belief system, “Kuru orthodoxy”) quickly expanded all over Northern India, and subsequently, across the Vindhya, to South India and later to S.E. Asia, up to Bali.

make him the culture hero of the East is as conspicuous as the R_{}˚ gvedic person Namin Sāpya as King of Videha at PB 25.10.17, where he is interestingly described as making a “pilgrimage” to Kuruks.etra (the holy land of the Veda and the home of Gotama Rāhūgan.a and Videgha Māthava who in ŚB are the prototypes of the eastward movement of Vedic orthopraxy).

159R. Hauschild, Über die frühesten Arier im Alten Orient, p. 55 on the migration of Videgha; A. Weber, Zwei Sagen aus dem Śatapatha-Brāhman.a über die Einwanderung und Verbreitung der Arier in Indien... Indische Studien 1,161 sqq.

160Unless one takes the ŚB legend literally and supposes a R_{}˚ gveda time move towards the east.

161To use H. Kulke’s term.

162Which nevertheless is attested: “The people move eastwards victoriously,” in the earlier YV text, KS 26.2:123.17, cf. Rau, Staat, p. 13.

163See Arrian, Indike 4.5 and cf. the commentary by O. v. Hinüber, in: Arrian, Der Alexanderzug. Indische Geschichte, hg. und übers. von G. Wirth u. O. v. Hinüber, München u. Zürich (Artemis) 1985, p. 1095; cf. also Author, Localisation.

164Note the two level settlements of small villages and larger market places of the Ochre Colored Pottery and Black and Red Ware cultures in the area.

165See above, notes 93–95.

This procedure is visible in the Bharata/Kaurava dynasty’s large scale collection of older and more recent religious texts: In all aspects of ritual, language and text collection, these texts tend to be more archaic than much of the inherited older texts and rites.^{167} On the other hand, the new dynasty was effective in re-shaping society^{168} and its structure by stratification into the four classes (varn.a), with an internal opposition between ārya and śūdra which effectively camouflaged the really existing social conflict between brahma-ks.atra and the rest, the vaiśya and śūdra; further, the Bhārata/Pāriks.ita dynasty was successful in reorganizing much of the traditional ritual and the texts concerned with it. (It must not be forgotten that public ritual included many of the functions of our modern administration, providing exchanges of goods, forging unity and underlining the power of the elite.)

The small tribal chieftainships of the R̊gvedic period with their shifting alliances and their history of constant warfare, though often not more than cattle rustling expeditions, were united in the single “large chiefdom” of the Kuru realm. With some justification, we may now call the great chief (rājā)^{169} of the Kurus “the Kuru king”. His power no longer depended simply on ritual relationships such as exchange of goods (vidatha)^{170} but on the extraction of tribute (bali)^{171} from an increasingly suppressed third estate (viś) and from dependent subtribes and weak neighbors; this was often camouflaged as ritual tribute, such as in the aśvamedha.

In view of the data presented in this paper, we are, I believe, entitled to call the Kuru realm the first state in India.^{172} To quote W. Rau, who has described the social and political conditions of the YV Sam .hitā and Brāhman.a period in such detail: “... the Indians of the Brāhman.a period lived in political organizations which, with good reasons, can be called states.”

It must be underlined, again, that the developments which brought about the the Kuru realm were lasting and not transient ones as those under the R̊gvedic Pūru or Bharata.

In effect, many of the changes in religion and society then carried out shape Indian society even TODAY.

166Ironically, if we use Srinivas’ term, — of people speaking Vedic Sanskrit.

167See above, notes 93–95.

168Note the Purus.a hymn R_{}˚ V 10.90, with its designation for the 4 classes (see above) and note that this is a late hymn indeed: it also has one of the first cases of loka- instead of uloka-.

169Note that even this title is traditional. While other areas of northern India use titles such as adhirāj, sarvarāj (see above) etc. the Kurus and Pañcālas retain the simple title of rājan. One may regard this as another “trick” of the Kuru king: he stresses that he is only a primus inter pares.

170Again, note that Pariks.it still uses this term at R_{}˚ VKh 5 — though he must have been the one who formed the early Kuru state.

171Rau, Staat, p. 104, cf. Rau, par. 24.3, 25, 40.2.

172Of course, barring the decipherment of the Indus seals which might point to a political and social organization that can be compared to Near Eastern states. See now W. A. Fairservis, The Harappan civilization and its Writing.

A Model for the Decipherment of the Indus Script. Delhi 1992.

ABBREVIATIONS:

AB  Aitareya Brāhman.a
ĀpMp  Āpastamba Mantrapāt.ha
ĀpŚS  Āpastamba Śrautasūtra
AV  Atharvaveda
AzI  K. Hoffmann, Aufsätze zur Indoiranistik
BĀU  Br̊hadāran.yakopanis.ad
Baudh  Baudhāyana
BĀUK  BĀU — Kān.va version
BĀUM  BĀU — Mādhyandina version
Bhār  Bhāradvāja
BhG  Bhagavadgītā
BŚS  Baudhāyana Śrautasūtra
DN  Dīghanikāya

EWA M. Mayrhofer, Etymologisches Wörterbuch des Altindoarischen, Heidelberg 1986–

GB  Gopatha Brāhman.a
Hir  Hiran.yakeśin
IE  Indo-European
IIJ  Indo-Iranian Journal
IIr  Indo-Iranian
JB  Jaiminīya Brāhman.a
JS  Jaiminīya Sam .hitā
JUB  Jaiminīya Upanis.adbrāhman.a
Kat.hUp  Kat.hopanis.ad
Kauś  Kauśika
KB  Kaus.ītaki Brāhman.a

KEWA M. Mayrhofer, Kurzgefasstes etymologisches Wörterbuch des Altindischen, Heidelberg 1956–1975

Kl.Schr.  Kleine Schriften
KpS  Kapis.t.hala-Kat.ha Sam .hitā
KS  Kat.ha Sam .hitā
KU  Kat.hopanis.ad
Kauth  Kauthuma
LŚS  Lāt.yāyana Śrautasūtra
MS  Maitrāyan.ī Sam .hitā
PB  Pañcavi ˙m̆śa Brāhman.a
PS  Paippalāda Sam .hitā

R̊V  R̊gveda (Sam .hitā)
R̊VKh  R̊gveda Khila
SV  Sāmaveda (Sam .hitā)
StII  Studien zur Indologie und Iranistik
TĀ  Taittirīya Āran.yaka
Taitt  Taittirīya
TB  Taittirīya Brāhman.a
TS  Taittirīya Sam .hitā
Vādh  Vādhūla
VādhB  Vādhūla Brāhman.a
VādhS  Vādhūla Sūtra
Vaikh  Vaikhānasa
VS  Vājasaneyi Sam .hitā
VSK  VS — Kan.va version
VSM  VS — Mādhyandina version
YV  Yajurveda (Sam .hitā)
Śāt.y  Śātyāyana
ŚB  Śatapatha Brāhman.a
ZDMG  Zeitschrift der deutschen morgenländischen Gesellschaft
ŚS  Śaunaka Sam .hitā
ŚŚS  Śāṅkhāyana Śrautasūtra